# P.808 Multilingual Speech Enhancement Testing: Approach and Results of URGENT 2025 Challenge

### v1

@@ Abstract
In speech quality estimation for speech enhancement (SE) systems, subjective listening tests so far are considered as the gold standard. This should be even more true considering the large influx of new generative or hybrid methods into the field, revealing issues of some objective metrics. Efforts such as the Interspeech 2025 URGENT Speech Enhancement Challenge also involving non-English datasets add the aspect of multilinguality to the testing procedure. In this paper, we provide a brief recap of the ITU-T P.808 crowdsourced subjective listening test method. A first novel contribution is our proposed process of localizing both text and audio components of Naderi and Cutler’s implementation of crowdsourced subjective absolute category rating (ACR) listening tests involving text-to-speech (TTS). Further, we provide surprising analyses of and insights into URGENT Challenge results, tackling the reliability of (P.808) ACR subjective testing as gold standard in the age of generative AI. Particularly, it seems that for generative SE methods, subjective (ACR MOS) and objective (DNSMOS, NISQA) reference-free metrics should be accompanied by objective phone fidelity metrics to reliably detect hallucinations. Finally, in the accepted version, we will release our localization scripts and methods for easy deployment for new multilingual speech enhancement subjective evaluations according to ITU-T P.808.
@@ 1 Introduction
  • Open source: Upon acceptance, we will provide our annotated scripts to the community for easy adoption, reproduction, and use for crowdsourced multilingual speech quality assessments. https://placeholder-url.git/name/P.808-Localization
@@ 4.2 Subjective and Objective Results
Table 1 shows the MOS values and select objective metrics across all participating models (including the baseline) on four languages which form subsets of $\mathcal{D}^{\mathrm{test}}_{\mathrm{blind}}$. We observe high(est) scores for MOS and intelligibility / phone fidelity metrics (ESTOI, LPS) for the English subset of $\mathcal{D}^{\mathrm{test}}_{\mathrm{blind}}$. This may be partially attributed to the imbalance of training data, which features an additional 802 hours of English speech besides the English speech included in the 1750 hours of multilingual data. The unseen language Japanese shows decreased subjective quality accompanied by a dramatic drop of phonetic fidelity (LPS $=0.53$), which is not reflected by the ESTOI metric. Interestingly, we observe the same rank order among languages for MOS and LPS. While the unseen Japanese language shows overall best reference-free metric results (DNSMOS 2nd rank, NISQA 1st rank), its overall MOS score is lowest among the four languages.
@@
We see that the discriminative model #1 scores highest in all reference-based objective metrics PESQ (significantly highest $2.61$), ESTOI ($0.82$), and LPS ($0.73$), however, falls behind the top MOS-scoring model #3 consistently in each investigated language.
The generative model #13 shows by far poorest overall intelligibility and phone fidelity (ESTOI $=0.53$, LPS $=0.58$) of all reported methods in Table 3. This is in significant contrast to the top scores of method #13 in the reference-free objective metrics DNSMOS ($3.13$) and NISQA ($3.79$). We can conclude from this example that DNSMOS and NISQA (dependent on the speech enhancement technology) may ignore the aspect of intelligibility and phone fidelity, whereas former literature states that speech quality includes intelligibility to some extent; both correlate until intelligibility is saturated [47, 48]. Chiang et al. [49] find that subjective quality and subjective intelligibility have a Pearson correlation coefficient exceeding even the correlation between subjective quality and PESQ, and between subjective quality and DNSMOS. Informal inspection of method #13 in DE and in EN confirms a high degree of hallucination. To summarize, we can state that reference-free objective metrics may very much ignore aspects such as intelligibility and phone fidelity (or hallucination).
It becomes even worse: Also the high MOS score ($3.34$) of method #13 is in surprising contrast to the low intelligibility and phone fidelity scores of method #13. We could state: ”It sounds like English [or any other language of the test], but it is not.” To the best of our knowledge, for the first time, the irritating observation is reported that (P.808) ACR listening tests (which are by nature also reference-free just as DNSMOS and NISQA) may simply ignore aspects such as intelligibility and phone fidelity (or hallucination). In the age of generative AI, can subjective ACR listening tests then still serve as gold standard? We tend to strongly recommend to ensure native listeners when judging generative speech enhancement methods, which is, however, difficult in the ITU-T P.808 crowdsourcing-based testing protocol. Our recommendation is that MOS results from P.808 ACR listening tests shall be accompanied by metrics that test for intelligibility (e.g., ESTOI), or for phone fidelity (e.g., LPS). This justifies the URGENT Challenges’ approach of a ranking obtained from average rankings in many objective metrics and MOS [5, 10]. High reference-free subjective or objective metrics coming along with a drop in ESTOI or particularly in LPS may signify a high degree in hallucinations.

### v2

@@ Abstract
In speech quality estimation for speech enhancement (SE) systems, subjective listening tests so far are considered as the gold standard. This should be even more true considering the large influx of new generative or hybrid methods into the field, revealing issues of some objective metrics. Efforts such as the Interspeech 2025 URGENT Speech Enhancement Challenge also involving non-English datasets add the aspect of multilinguality to the testing procedure. In this paper, we provide a brief recap of the ITU-T P.808 crowdsourced subjective listening test method. A first novel contribution is our proposed process of localizing both text and audio components of Naderi and Cutler’s implementation of crowdsourced subjective absolute category rating (ACR) listening tests involving text-to-speech (TTS). Further, we provide surprising analyses of and insights into URGENT Challenge results, tackling the reliability of (P.808) ACR subjective testing as gold standard in the age of generative AI. Particularly, it seems that for generative SE methods, subjective (ACR MOS) and objective (DNSMOS, NISQA) reference-free metrics should be accompanied by objective phone fidelity metrics to reliably detect hallucinations. Finally, we will soon release our localization scripts and methods for easy deployment for new multilingual speech enhancement subjective evaluations according to ITU-T P.808.
@@ 1 Introduction
  • Open source: Soon, we will provide our annotated scripts to the community for easy adoption, reproduction, and use for crowdsourced multilingual speech quality assessments. https://placeholder-url.git/name/P.808-Localization
@@ 4.2 Subjective and Objective Results
Table 1 shows the MOS values and select objective metrics across all participating models (including the baseline) on four languages which form subsets of $\mathcal{D}^{\mathrm{test}}_{\mathrm{blind}}$. We observe high(est) scores for MOS and intelligibility / phone fidelity metrics (ESTOI, LPS) for the English subset of $\mathcal{D}^{\mathrm{test}}_{\mathrm{blind}}$. This may be partially attributed to the imbalance of training data, which features an additional 802 hours of English speech besides the English speech included in the 1750 hours of multilingual data. The unseen language Japanese shows decreased subjective quality accompanied by a dramatic drop of phonetic fidelity (LPS =0.53=0.53= 0.53), which is not reflected by the ESTOI metric. Interestingly, we observe the same rank order among languages for MOS and LPS. While the unseen Japanese language shows overall best reference-free metric results (DNSMOS 2nd rank, NISQA 1st rank), its overall MOS score is lowest among the four languages.
@@
We see that the discriminative model #1 scores highest in all reference-based objective metrics PESQ (significantly highest 2.612.612.61), ESTOI (0.820.820.82), and LPS (0.730.730.73), however, falls behind the top MOS-scoring model #3 consistently in each investigated language.
The generative model #13 shows by far poorest overall intelligibility and phone fidelity (ESTOI =0.53=0.53= 0.53, LPS =0.58=0.58= 0.58) of all reported methods in Table 3. This is in significant contrast to the top scores of method #13 in the reference-free objective metrics DNSMOS (3.133.133.13) and NISQA (3.793.793.79). We can conclude from this example that DNSMOS and NISQA (dependent on the speech enhancement technology) may ignore the aspect of intelligibility and phone fidelity, whereas former literature states that speech quality includes intelligibility to some extent; both correlate until intelligibility is saturated [47, 48]. Chiang et al. [49] find that subjective quality and subjective intelligibility have a Pearson correlation coefficient exceeding even the correlation between subjective quality and PESQ, and between subjective quality and DNSMOS. Informal inspection of method #13 in DE and in EN confirms a high degree of hallucination. To summarize, we can state that reference-free objective metrics may very much ignore aspects such as intelligibility and phone fidelity (or hallucination).
It becomes even worse: Also the high MOS score (3.343.343.34) of method #13 is in surprising contrast to the low intelligibility and phone fidelity scores of method #13. We could state: ”It sounds like English [or any other language of the test], but it is not.” To the best of our knowledge, for the first time, the irritating observation is reported that (P.808) ACR listening tests (which are by nature also reference-free just as DNSMOS and NISQA) may simply ignore aspects such as intelligibility and phone fidelity (or hallucination). In the age of generative AI, can subjective ACR listening tests then still serve as gold standard? We tend to strongly recommend to ensure native listeners when judging generative speech enhancement methods, which is, however, difficult in the ITU-T P.808 crowdsourcing-based testing protocol. Our recommendation is that MOS results from P.808 ACR listening tests shall be accompanied by metrics that test for intelligibility (e.g., ESTOI), or for phone fidelity (e.g., LPS). This justifies the URGENT Challenges’ approach of a ranking obtained from average rankings in many objective metrics and MOS [5, 10]. High reference-free subjective or objective metrics coming along with a drop in ESTOI or particularly in LPS may signify a high degree in hallucinations.